[Brahms-l] Shift report 20050224 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Feb 24 2005 - 23:54:34 EST
Shifter leader was HHD
Shifters were 


1. As soon as the beam was up and running, things went smoothly until the RHIC needed repair at 2330. The expected downtime was 90 minutes.

2. Runs taken: (tr3 and tr6 off as angle is 2.3)
   FS@2.3  A2442: 13472 (265k)
           A3450: 13473 (???), 13374 (252k), 13475 (61k)

Non-TPC pedestal run: 13476
TPC pedestal run 13477

3. The run server crashed during run 13473. As a result the run could not be ended in the e-log, but I think the data is there. Perhaps somebody could fix it? Apart from this there were no problems (the TPCs were especially cooperative).  

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Feb 24 23:55:54 2005

This archive was generated by hypermail 2.1.8 : Thu Feb 24 2005 - 23:56:07 EST